25 June 2000  Page 1 of 1  J3/00-242

Subject: Issue 237 From: Van Snyder

1 Introduction

Issue 237 is concerned in part with the definition of NEAREST rounding during input/output. It argues that the definition of NEAREST is redundant.

2 Edits

Edits refer to 00-007r2. Page and line numbers are displayed in the margin. Absent other instructions, a page and line number or line number range implies all of the indicated text is to be replaced by immediately following text, while a page and line number followed by + indicates that immediately following text is to be inserted after the indicated line. Remarks for the editor are noted in the margin, or appear between [ and ] in the text.

[Editor: Delete “shall be ... values used.”] 224:29-30

[Editor: Delete the first paragraph of issue 237.] 225:3-97